OpenStudy (shesolitt):

What is the slope of the line y = 3? Undefined 0 << not sure if its B −3 3 << not sure if its D

OpenStudy (math&ing001):

Remember when we have y = mx + b, m is the slope Likewise y = 0*x + 3 What do you think ?

OpenStudy (shesolitt):

i think its 3 @math&ing001

OpenStudy (math&ing001):

y = mx + b m is the slope and b is the y-intercept y = 0*x + 3 0 is the slope and 3 is the y-intercept

OpenStudy (shesolitt):

thank you i got a 100% @math&ing001

OpenStudy (math&ing001):

Welcome !